Use synthetic division to find the quotient Q(x) and remainder R when
P(x)= -x^3 +24x -5 is divided by x+5 . Specifically, complete the synthetic division table below, and write your answer in the form Q(x) +R/(x+5).

-5/ -1, 0, 24, -5
?, ?, ?
______________________
?, ?, ?, ?

-x^3 +24x -5/ (x+5) = ? + ?/(x+5)

Answer by nerdybill(7384)

Use synthetic division to find the quotient Q(x) and remainder R when
P(x)= -x^3 +24x -5 is divided by x+5 . Specifically, complete the synthetic division table below, and write your answer in the form Q(x) +R/(x+5).

-5/ -1, 0, 24, -5
5,-25, 5
______________________
-1, 5, -1, 0

-x^3 +24x -5/ (x+5) = -x^2+5x-1+0/(x+5)
OR, simply
-x^3 +24x -5/ (x+5) = -x^2+5x-1